XYZ estimates that if it issues new common stock, the flotation cost will be 15%. XYZ incorporates the flotation costs into the DCF approach. What is the estimated cost of newly issued common stock, considering the flotation cost? Please show work in Excel.

7. (12.5 points) XYZ estimates that if it issues new common stock, the flotation cost will be 15%. XYZ incorporates the flotation costs into the DCF approach. What is the estimated cost of newly issued common stock, considering the flotation cost? 15% $42.50 % Flotation cost Net proceeds after flotation Cost of Newly Issued Common Stock
Answer #1

cost of newly issued comon stock = 4.19(1.05)/50(1-.015)+.05

= 4.40/49.25

=.137 or 13.7%

alternative = 4.40/42.5+.05=.154 or 15.4%

between 13.7% -----15.4% will be the new cost